Give me two examples with solutions of each types of discontinuities

1
Expert's answer
2022-02-23T10:05:12-0500

Solution:

Removable discontinuity:

Example 1:

"f(x)=\frac{x^{2}-5 x+4}{x-4}"

Find the factors of the numerator and the denominator. The denominator cannot be factored further, but we can factor the numerator.

"\begin{gathered}

=x^{2}-5 x+4 \\

=x^{2}-4 x-x+4 \\

=x(x-4)-1(x-4) \\

=(x-1)(x-4) \\

f(x)=\frac{(x-1)(x-4)}{x-4}

\end{gathered}"

Now, we have identified the common factor which is x-4 . We need to set it equal to zero to get the removable discontinuity.

"\begin{aligned}

&x-4=0 \\

&x=4

\end{aligned}"

Hence, the removable discontinuity of the function is at the point x=4 .

Example 2:

"f(x)=\frac{x^{2}-7 x+12}{x-4}"

Find the factors of the numerator and the denominator. The denominator cannot be factored further, but we can factor the numerator.

"\begin{gathered}

=x^{2}-7 x+12 \\

=x^{2}-3 x-4x+12 \\

=x(x-3)-4(x-3) \\

=(x-3)(x-4) \\

f(x)=\frac{(x-3)(x-4)}{x-4}

\end{gathered}"

Now, we have identified the common factor which is x-4 . We need to set it equal to zero to get the removable discontinuity.

"\begin{aligned}

&x-4=0 \\

&x=4

\end{aligned}"

Hence, the removable discontinuity of the function is at the point x=4 .

Non-removable discontinuity:

Example 1:

Find the points of discontinuity of the "f(x)=\frac{1}{2 \sin x-1}" .

Solution: Given that "f(x)=\frac{1}{2 \sin x-1}" .

 f(x) is discontinuous when

"2 \sin x-1=0"

"\operatorname{Sin} x=1 / 2



\\x=\sin ^{-1}(1 / 2)



\\x=\pi / 6 \text { (or) } 5 \pi / 6"

The general solution is "2 n \pi \pm \frac{\pi}{6} (or) 2 n \pi \pm \frac{5 \pi}{6}, \mathrm{n} \in \mathrm{z} ."

Example 2:

Find the points of discontinuity of the "f(x)=\frac{1}{2x+1}" .

Solution: Given that "f(x)=\frac{1}{2x+1}" .

 f(x) is discontinuous when

"2 x+1=0"

"2x=-1
\\x=\dfrac{-1}2"

f(x) is discontinuous at "x=\dfrac{-1}2"